# TrialSynth: Generation of Synthetic Sequential Clinical Trial Data

### v1

@@ 4 Experiments
LC1 0.689±0.105 0.563±0.053 0.504±0.066 0.508±0.122 0.557±0.055 0.457±0.079 0.572±0.051 0.709±0.049
BC1 0.678±0.078 0.617±0.036 0.573±0.043 0.550±0.046 0.630±0.045 0.461±0.184 0.646±0.037 0.665±0.045
CC 0.637±0.140 0.481±0.092 0.567±0.096 0.448±0.023 0.583±0.098 0.446±0.02 0.622±0.016 0.653±0.019
BC2 0.660±0.128 0.535±0.073 0.523±0.074 0.523±0.11 0.513±0.078 0.503±0.075 0.548±0.024 0.594±0.068
BC3 0.632±0.072 0.454±0.039 0.463±0.039 0.493±0.013 0.503±0.043 0.535±0.183 0.590±0.050 0.634±0.032
VTE 0.640±0.038 0.490±0.019 0.549±0.022 0.508±0.113 0.531±0.021 0.485±0.066 0.568±0.018 0.625±0.020
LC2 0.738±0.149 0.563±0.097 0.507±0.087 0.573±0.118 0.574±0.096 0.534±0.078 0.689±0.084 0.755±0.059
4.1 Datasets
We evaluated our models on 7 real-world clinical trial outcome datasets obtained from Project Data Sphere [^3] https://data.projectdatasphere.org/projectdatasphere/html/access (Green et al., 2015; Fu et al., 2023; Chang et al., 2019; Chen et al., 2024b, c). Specifically, we chose the trials as outlined in Table 1. These datasets have shown to be effective evaluation datasets for tabular prediction (Wang & Sun, 2022b; Wang et al., 2023a) and digital twin generation (Das et al., 2023; Wang et al., 2024). Specifically, we use LC1 (Niell et al., 2005), BC1 (Baldwin et al., 2012), CC (Alberts et al., 2012), BC2 (Fernández-Cuesta et al., 2012), BC3 (WU et al., 2024), VTE (Agnelli et al., 2012), LC2 (Daniel et al., 2021; Fu et al., 2024; Zhang et al., 2021). A full description of the data is shown in Table 1. Each dataset contains events and the times at which they occur, e.g., medications and procedures, as well as some adverse events like vomiting etc. We use these datasets to predict if the subject experiences the death event, which is an external label. Note that TrialSynth does not require a fixed patient event sequence length.
4.2 Baseline Methods
As stated in the related work, existing EHR methods and synthetic patient generation methods are not applicable to our specific task and dataset due to dataset size and lack of support for timestamp generation; therefore, we primarily compare against general sequential data generation methods. We compared the following 7 models: First, the LSTM VAE is our implementation of a variational autoencoder using LSTM, designed to predict event types and timestamps for each event as a categorical classification and regression task, respectively. PARSynthesizer from is SDV, based on a conditional probabilistic auto-regressive (CPAR) model, is specifically tailored for synthesizing sequential event data and stands out due to its unique focus and accessible codebase. TabDDPM is a state-of-the-art tabular synthesizer using diffusion models, enhanced by adding time as a numerical column for our purposes. Despite not being explicitly designed for sequential data, it surpasses previous models like CTGAN in synthetic tabular data generation. Lastly, HALO, a hierarchical autoregressive language model, excels in synthesizing Electronic Health Records (EHR) but struggles with clinical trial datasets due to the limited size of the training data, highlighting the challenges in this domain.
TrialSynth (Events Unknown) is the VAE + Multivariate Hawkes Process that is trained without any assumptions. At training time, the task is to predict the next event and timestamp given a history of observations. At inference time, we perform autoregressive decoding, adding the last prediction to the history until a predicted cutoff is reached. TrialSynth (Events Known) assumes that one knows which specific events occur for the Hawkes Model (but not the timestamps or the number of occurrences). Similar to TrialSynth (Events Unknown), training is performed by predicting the next event and timestamp given history, and autoregressive decoding is done at inference time until a predicted cutoff.
4.3 Utility Evaluation
Downstream Classification ROCAUC: It is vital that synthetic data perform similarly to real-world data; therefore, we evaluate the utility (ROCAUC) of the generated synthetic data by performing binary classification of death events in all 7 clinical trials. We choose ROCAUC since it has been used for similar tasks in the pastDas et al. (2023). Additionally, ROC AUC is sensitive to class imbalance in the sense that when there is a minority class, one typically defines this as the positive class and it will have a strong impact on the AUC value. This is desirable behavior and is what we were looking to evaluate in our application.
@@ 4.3 Utility Evaluation
We see that synthetic data generated by TrialSynth variants generally perform the best in terms of downstream death event classification performance, where TrialSynth (Events Unknown) outperforms the next best model (in 4/7 datasets and is within 1 standard deviation with the rest of the datasets). Furthermore, TrialSynth (Events Known) significantly outperforms other baselines, due to the additional input information.
@@ 4.4 Privacy evaluations
LC2 1.00 1.00 0.91 1.00 0.27 0.72 0.25
ML Inference Score: This can also be thought of as an adversarial Model Attack Theodorou et al. (2023). Another main concern is the privacy of the synthetic data, to prevent any data or information leakage. To address this, we calculate the performance of predicting whether a generated sequence is real vs synthetic via an LSTM binary classification (Patki et al., 2016) (similar to an adversarial model). The real subjects are labeled with “0” and the synthetic subjects are labelled with “1”. Results are shown in Table 3, and we see that TrialSynth variants perform closest to the optimal 0.5 ROCAUC ideal score. One thing to note is that a perfect copy of the original data would result in a 0.5 score, so we have the following metric to measure the opposite scenario. Furthermore, we see a continued trend of TrialSynth (Events Known) generally outperforming other baseline methods, illustrating the importance of giving the model more information in this data-scarce setting.
Distance to Closest Record (DCR) Score: Second, we follow the evaluation metrics per TabDDPM (Kotelnikov et al., 2023). That is, we compare the feature vectors of the real vs synthetic data and measure how far the synthetic data is from the original. The higher this distance is, the more different the generated data is from the original data, and thus the more private it is. A completely different version of the data would obtain the highest distance but could result in bad performance in the downstream LSTM classification performance or a high ML Inference score (close to 1). We calculate this by featurizing the event time predictions in terms of (count, means, and standard deviations). Then, we normalize and obtain the L2 distance between a generated subject and the closest real subject. Table 5 shows this result. Notice that TrialSynth variants generally obtain quite low scores on this metric. TabDDPM and PAR also generate data closer to the original data compared to LSTM VAE. We note the privacy-fidelity trade-off, as LSTM VAE generates data that is further away from the original, but yields worse utility (Table 2).
Dataset Attack Per HALO Theodorou et al. (2023), we also added a Dataset Attack evaluation, where we label the real records with the lowest distance (computed by featuring event times into mean, std, counts) to the closest record in the synthetic dataset as 1. It tests the ability of the synthetic dataset to prevent an attacker from inferring whether a real record was used in the training dataset. On real training data, we compare if the closest record is a real record from the training set or a synthetic record. Ideally, we also want this accuracy to be 0.5.
From Table 5, we see that TrialSynth generally performs the best, even beating out HALO and TabDDPM.
@@ 4.5 Utility / Privacy Trade-off
Here, we visualize 3 utility/privacy trade-off plots that is inherent to any synthetic data generation task. Each metric is normalized for ease of visualization so that the maximum achieved metric is set as the tip of the triangle by dividing by the max. For ML Inference Privacy (where 0.5 is the ideal value), we first take the absolute value of the difference (i.e. $x=|x-0.5|$), and then divide by the max as before. The results are shown in Figure 7. It is important to note that while a max score of 1 would be ideal for all models, in practice, we see a clear trade-off as the best-performing Distance to Closest Record model. VAE LSTM and PAR perform worse on the downstream ROCAC metric compared to more faithful (and higher utility in terms of downstream ROCAUC) models such as TabDDPM and TrialSynth. This is because the generated sequences are of worse fidelity, being too different from the original. TrialSynth variants also generally have good ML Inference Privacy, which is to be expected as those models generate data that is similar to the original, which would allow for (1) better performance on the held-out test set for ROCAUC and (2) being harder to distinguish from original data.
@@ 5 Discussion
We introduce TrialSynth, a novel model that combines Variational Autoencoder and Hawkes Process techniques. While we have evaluated TrialSynth on clinical trial patient data, the methodology is quite general and can be applied to other forms of sequential data, such as financial or social media data. This model excels at generating sequential event data with precise timestamps, and we demonstrate its superior performance compared to the traditional sequential models (LSTM and PAR) models in handling tabular data for downstream machine learning tasks.
Specifically, we demonstrate that TrialSynth outperforms existing methods in terms of data utility, enabling the generation of highly authentic event sequences across multiple real-world sequential event datasets. Empirical experiments indicate that providing the model with additional information, such as event index (Events Known) or event length, leads to significant improvements in the synthetic data quality. Finally, we believe that a sweet spot is reached by allowing the model to know the event index–as it provides a significant downstream classification boost while maintaining a low ML inference score, and is a common assumption when generating specific patients.
@@ A.1.1 Societal Impact
A.2 TrialSynth Details
Neural Hawkes Processes are formulated as follows. We are given a set of $L$ observations of the form (time $t_{j}$, event_type $k_{j}$). $S=\{(t_{1},k_{1}),\dots,(t_{j},k_{j}),\dots,(t_{L},k_{L})\}$ Each time $t_{j}\in\mathbb{R}^{+}\bigcup\{0\}$ and is sorted such that $t_{j}<t_{j+1}$. Each event $k_{j}\in\{1,\dots,K\}$. The traditional Hawkes Process assumption that events only have a positive, decaying influence on future events is not realistic in practice, as there exist examples where an occurrence of an event lowers the probability of a future event (e.g., medication reduces the probability of adverse events). Therefore, the Neural Hawkes Process (Mei & Eisner, 2017) was proposed to generalize the traditional Hawkes Process. The following derivations follow (Zuo et al., 2020).
@@ A.2 TrialSynth Details
For training, we want to maximize the likelihood of the observed sequence $\{(t_{1},k_{1}),\dots,(t_{L},k_{L})\}$. The log-likelihood function is given by [^4] The proof is shown in (Mei & Eisner, 2017)
@@
Finally, since the gradient of the log-likelihood function has an intractable integral, one may obtain an unbiased estimate by performing Monte Carlo sampling (Robert et al., 1999).
@@
For the VAE loss, we want to minimize the Kullback–Leibler divergence between $q_{\phi}(z|x)$ and $p_{\theta}(z|x)$, which in practice leads to maximizing the evidence lower bound (ELBO) for training along with the likelihood of $x$ (Kingma & Welling, 2013).
@@
A.3 Ethics and Reproducibility
Transformer Hawkes (Zuo et al., 2020) is open source and can be found at https://github.com/SimiaoZuo/Transformer-Hawkes-Process. Training on an NVIDIA GeForce RTX 3090 takes around 12 hrs to run the full model. The code will be made public and open source on GitHub. for the camera-ready version. All datasets were obtained from Project Data Sphere (Green et al., 2015) with permission via a research data access request form. The links are as follows:
@@ A.3 Ethics and Reproducibility
    NCT00003299 (Niell et al., 2005): A Randomized Phase III Study Comparing Etoposide and Cisplatin With Etoposide, Cisplatin and Paclitaxel in Patients With Extensive Small Cell Lung Cancer. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/261
@@
    NCT00041119 (Baldwin et al., 2012): Cyclophosphamide And Doxorubicin (CA) (4 VS 6 Cycles) Versus Paclitaxel (4 VS 6 Cycles) As Adjuvant Therapy For Breast Cancer in Women With 0-3 Positive Axillary Lymph Nodes:A 2X2 Factorial Phase III Randomized Study. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/486
@@
    NCT00079274 (Alberts et al., 2012): A Randomized Phase III Trial of Oxaliplatin (OXAL) Plus 5-Fluorouracil (5-FU)/Leucovorin (CF) With or Without Cetuximab (C225) After Curative Resection for Patients With Stage III Colon Cancer. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/407
@@
    NCT00174655 (Fernández-Cuesta et al., 2012): An Intergroup Phase III Trial to Evaluate the Activity of Docetaxel, Given Either Sequentially or in Combination With Doxorubicin, Followed by CMF, in Comparison to Doxorubicin Alone or in Combination With Cyclophosphamide, Followed by CMF, in the Adjuvant Treatment of Node-positive Breast Cancer Patients. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/127
@@
    NCT00312208 (Mackey et al., 2016): A Multicenter Phase III Randomized Trial Comparing Docetaxel in Combination With Doxorubicin and Cyclophosphamide Versus Doxorubicin and Cyclophosphamide Followed by Docetaxel as Adjuvant Treatment of Operable Breast Cancer HER2neu Negative Patients With Positive Axillary Lymph Nodes. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/118
@@
    NCT00694382 (Agnelli et al., 2012): A Multinational, Randomized, Double-Blind, Placebo-controlled Study to Evaluate the Efficacy and Safety of AVE5026 in the Prevention of Venous Thromboembolism (VTE) in Cancer Patients at High Risk for VTE and Who Are Undergoing Chemotherapy. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/119
@@
    NCT03041311 (Daniel et al., 2021): Phase 2 Study of Carboplatin, Etoposide, and Atezolizumab With or Without Trilaciclib in Patients With Untreated Extensive-Stage Small Cell Lung Cancer (SCLC). Available at https://data.projectdatasphere.org/projectdatasphere/html/content/435
A.4 Baselines

### v2

@@ 4.3 Utility Evaluation
LC1 0.689±0.105 0.563±0.053 0.504±0.066 0.508±0.122 0.557±0.055 0.457±0.079 0.672±0.061 0.709±0.049
BC1 0.678±0.078 0.617±0.036 0.573±0.043 0.550±0.046 0.630±0.045 0.461±0.184 0.651±0.046 0.665±0.045
CC 0.657±0.140 0.481±0.092 0.567±0.096 0.448±0.023 0.583±0.098 0.446±0.02 0.652±0.015 0.653±0.019
BC2 0.660±0.128 0.535±0.073 0.523±0.074 0.523±0.11 0.513±0.078 0.503±0.075 0.599±0.042 0.594±0.068
BC3 0.632±0.072 0.454±0.039 0.463±0.039 0.493±0.013 0.503±0.043 0.535±0.183 0.620±0.038 0.634±0.032
VTE 0.640±0.038 0.490±0.019 0.549±0.022 0.508±0.113 0.531±0.021 0.485±0.066 0.618±0.024 0.625±0.020
LC2 0.738±0.149 0.563±0.097 0.507±0.087 0.573±0.118 0.574±0.096 0.534±0.078 0.729±0.044 0.755±0.059
Downstream Classification ROCAUC: It is vital that synthetic data perform similarly to real-world data; therefore, we evaluate the utility (ROCAUC) of the generated synthetic data by performing binary classification of death events in all 7 clinical trials. We choose ROCAUC since it has been used for similar tasks in the past[12]. Additionally, ROC AUC is sensitive to class imbalance in the sense that when there is a minority class, one typically defines this as the positive class and it will have a strong impact on the AUC value. This is desirable behavior and is what we look to evaluate in our application.
@@
We see that synthetic data generated by TrialSynth variants generally perform the best in terms of downstream death event classification performance, where TrialSynth (Events Unknown) outperforms the next best model (in 4/7 datasets and is within 1 standard deviation with the rest of the datasets). Furthermore, TrialSynth (Events Known) significantly outperforms other baselines, due to the additional input information. Still, TrialSynth (Events Unknown) also performs admirably, being on par but slightly less performant than TrialSynth (Events Known).
@@ 4.4 Privacy evaluations
ML Inference Score: This can also be thought of as an adversarial Model Attack [36]. Another main concern is the privacy of the synthetic data, to prevent any data or information leakage. To address this, we calculate the performance of predicting whether a generated sequence is real vs synthetic via an LSTM binary classification [33] (similar to an adversarial model). The real subjects are labeled with “0” and the synthetic subjects are labelled with “1”. Results are shown in Table 3, and we see that TrialSynth variants perform closest to the optimal 0.5 ROCAUC ideal score. One thing to note is that a perfect copy of the original data would result in a 0.5 score, so we have the following metric to measure the opposite scenario. Furthermore, we see a continued trend of both forms of TrialSynth generally outperforming other baseline methods, illustrating the importance of giving the model more information in this data-scarce setting.
@@
LC2 1.00 1.00 0.91 1.00 0.27 0.62 0.25
Distance to Closest Record (DCR) Score: Second, we follow the evaluation metrics per TabDDPM [25]. That is, we compare the feature vectors of the real vs synthetic data and measure how far the synthetic data is from the original. The higher this distance is, the more different the generated data is from the original data, and thus the more private it is. A completely different version of the data would obtain the highest distance but could result in bad performance in the downstream LSTM classification performance or a high ML Inference score (close to 1). We calculate this by featurizing the event time predictions in terms of (count, means, and standard deviations). Then, we normalize and obtain the L2 distance between a generated subject and the closest real subject. Table 5 shows this result. Notice that TrialSynth variants generally obtain quite low scores on this metric. TabDDPM and PAR also generate data closer to the original data compared to LSTM VAE. We note the privacy-fidelity trade-off, as LSTM VAE generates data that is further away from the original, but yields worse utility (Table 2).
Dataset Attack We evaluate a Dataset Attack scenario as per HALO [36], where we label the real records with the lowest distance (computed by featuring event times into mean, std, counts) to the closest record in the synthetic dataset as 1. It tests the ability of the synthetic dataset to prevent an attacker from inferring whether a real record was used in the training dataset. On real training data, we compare if the closest record is a real record from the training set or a synthetic record. Ideally, we also want this accuracy to be 0.5. From Table 5, we see that TrialSynth generally performs the best, even beating out HALO and TabDDPM.
@@ 4.5 Utility / Privacy Trade-off
Figure 4: 2 Privacy-Utility Tradeoff examples in TrialSynth: Performance of distance to closest record (DCR) (red) and downstream ROC (blue) metrics at varying levels of VAE sampling variance (from 0.1 to 4), represented as the “Var Multiplier.”
In TrialSynth, the privacy-utility tradeoff is governed by the variance applied to the VAE sampling process (Figure 4 and Figure 7). Increasing the variance in VAE sampling introduces more diversity into the synthetic data, enhancing privacy by making it harder to trace back to original data points. However, as the Var Multiplier rises, the quality of utility metrics such as downstream ROC tends to decrease, reflecting a drop in predictive accuracy and utility for downstream tasks. Conversely, metrics like DCR may rise, indicating a more extensive departure from the original dataset. A unique advantage of TrialSynth is its capacity to provide direct control over the tradeoff between fidelity and privacy through the adjustment of VAE sampling variance. By tuning this "Var Multiplier," researchers can precisely regulate how closely the synthetic data resembles the original dataset. Lower variance settings yield data with higher fidelity, making it more useful for predictive analyses and downstream clinical tasks, while higher variance introduces greater diversity, enhancing privacy protections by reducing the likelihood of re-identifying individual patients.
@@ 5 Discussion
The study presents TrialSynth, an innovative model that combines Variational Autoencoders (VAE) with Hawkes Processes (HP) to generate realistic synthetic sequential clinical trial data. Designed to address the challenges of small patient populations and the need for detailed time-event sequences, TrialSynth effectively captures both the timing and type of clinical events with high fidelity. Compared to existing methods, it outperforms in preserving data utility for downstream tasks while maintaining robust privacy protections, making it difficult to distinguish synthetic data from real data. Specifically, we demonstrate that TrialSynth outperforms existing methods in terms of data utility, enabling the generation of highly authentic event sequences across multiple real-world sequential event datasets. Empirical experiments indicate that providing the model with additional information, such as event index (Events Known) or event length, leads to significant improvements in the synthetic data quality. Finally, we believe that a sweet spot is reached by allowing the model to know the event index–as it provides a significant downstream classification boost while maintaining a low ML inference score, and is a common assumption when generating specific patients. We note that relaxing this assumption still yields competitive performance. Overall, TrialSynth offers a powerful solution for synthetic data generation in healthcare, balancing patient privacy with data authenticity, and shows promise for broader applications in clinical trial design and other healthcare domains that demand high-quality, secure synthetic datasets.
@@ A.2 Societal Impact
A.3 TrialSynth Details
Neural Hawkes Processes are formulated as follows. We are given a set of $L$ observations of the form (time $t_{j}$, event_type $k_{j}$). $S=\{(t_{1},k_{1}),\dots,(t_{j},k_{j}),\dots,(t_{L},k_{L})\}$ Each time $t_{j}\in\mathbb{R}^{+}\bigcup\{0\}$ and is sorted such that $t_{j}<t_{j+1}$. Each event $k_{j}\in\{1,\dots,K\}$. The traditional Hawkes Process assumption that events only have a positive, decaying influence on future events is not realistic in practice, as there exist examples where an occurrence of an event lowers the probability of a future event (e.g., medication reduces the probability of adverse events). Therefore, the Neural Hawkes Process [30] was proposed to generalize the traditional Hawkes Process. The following derivations follow [51].
@@ A.3 TrialSynth Details
For training, we want to maximize the likelihood of the observed sequence $\{(t_{1},k_{1}),\dots,(t_{L},k_{L})\}$. The log-likelihood function is given by [^4] The proof is shown in [30]
@@
Finally, since the gradient of the log-likelihood function has an intractable integral, one may obtain an unbiased estimate by performing Monte Carlo sampling [34].
@@
For the VAE loss, we want to minimize the Kullback–Leibler divergence between $q_{\phi}(z|x)$ and $p_{\theta}(z|x)$, which in practice leads to maximizing the evidence lower bound (ELBO) for training along with the likelihood of $x$ [24].
@@
A.4 Ethics and Reproducibility
Transformer Hawkes [51] is open source and can be found at https://github.com/SimiaoZuo/Transformer-Hawkes-Process. Training on an NVIDIA GeForce RTX 3090 takes around 12 hrs to run the full model. The code will be made public and open source on GitHub. for the camera-ready version. All datasets were obtained from Project Data Sphere [19] with permission via a research data access request form. The links are as follows:
@@ A.4 Ethics and Reproducibility
    NCT00003299 [31]: A Randomized Phase III Study Comparing Etoposide and Cisplatin With Etoposide, Cisplatin and Paclitaxel in Patients With Extensive Small Cell Lung Cancer. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/261
@@
    NCT00041119 [3]: Cyclophosphamide And Doxorubicin (CA) (4 VS 6 Cycles) Versus Paclitaxel (4 VS 6 Cycles) As Adjuvant Therapy For Breast Cancer in Women With 0-3 Positive Axillary Lymph Nodes:A 2X2 Factorial Phase III Randomized Study. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/486
@@
    NCT00079274 [2]: A Randomized Phase III Trial of Oxaliplatin (OXAL) Plus 5-Fluorouracil (5-FU)/Leucovorin (CF) With or Without Cetuximab (C225) After Curative Resection for Patients With Stage III Colon Cancer. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/407
@@
    NCT00174655 [14]: An Intergroup Phase III Trial to Evaluate the Activity of Docetaxel, Given Either Sequentially or in Combination With Doxorubicin, Followed by CMF, in Comparison to Doxorubicin Alone or in Combination With Cyclophosphamide, Followed by CMF, in the Adjuvant Treatment of Node-positive Breast Cancer Patients. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/127
@@
    NCT00312208 [29]: A Multicenter Phase III Randomized Trial Comparing Docetaxel in Combination With Doxorubicin and Cyclophosphamide Versus Doxorubicin and Cyclophosphamide Followed by Docetaxel as Adjuvant Treatment of Operable Breast Cancer HER2neu Negative Patients With Positive Axillary Lymph Nodes. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/118
@@
    NCT00694382 [1]: A Multinational, Randomized, Double-Blind, Placebo-controlled Study to Evaluate the Efficacy and Safety of AVE5026 in the Prevention of Venous Thromboembolism (VTE) in Cancer Patients at High Risk for VTE and Who Are Undergoing Chemotherapy. Available at https://data.projectdatasphere.org/projectdatasphere/html/content/119
@@
    NCT03041311 [11]: Phase 2 Study of Carboplatin, Etoposide, and Atezolizumab With or Without Trilaciclib in Patients With Untreated Extensive-Stage Small Cell Lung Cancer (SCLC). Available at https://data.projectdatasphere.org/projectdatasphere/html/content/435
A.5 Baselines